an object is placed 5 feet from the center of a horizontally rotating platform. the coefficient of friction is 0.2. the object will begin to slide off when platform speed is nearest to


1
Expert's answer
2021-06-28T17:10:02-0400

Gives

"\mu=0.2;R=5feet=1.524m"

"V=\sqrt{\mu R g}"

"V=\sqrt{0.2\times\times1.524\times9.8}"

"V=\sqrt{2.98704}" m/sec


"V=1.73m\/sec"